OpenStudy (anonymous):

ok x/x-4 / x/4-x so is the answer -x?

myininaya (myininaya):

Is that \[\frac{x}{x-4} \div \frac{x}{4-x} ?\]

myininaya (myininaya):

Hint: \[\frac{a}{b} \div \frac{c}{d}=\frac{a}{b} \cdot \frac{d}{c}\]